OpenStudy (anonymous):

3 cups are to 1 can as 9 cups are to cans.

OpenStudy (anonymous):

3 cans

OpenStudy (anonymous):

3 times 3 equals 9 so 3 cans

OpenStudy (anonymous):

\[\frac{ 3cups }{ 1can }=\frac{ 9cups }{ xcans }\] Find what is the value that multiplied times 3 = 9. Then use it on the bottom to find the number of cans.

OpenStudy (anonymous):

is it 3 cans amdkik me at sosakiing follow on instagram grandpark_kid

OpenStudy (kkutie7):

\[9cups*\frac{ 1can }{ 3cups }\] I always write things like this because it make its easy to see the next part. Now you want to cross out things that are opposite of each other. here we can see that we have one cups on the top and one on the bottom. Multiply. \[\frac{9cups*1can}{3cups}\] See how you can cross out the cups? you should be left with the unit cans now \[\frac{9cans}{3}\] divide to get 3cans =)